a catalog retailer is preparing to release a new catalog. they initially tested the response to the catalog in a subset of their customers. in the initial test, they reached out to 30,000 customers and sent them a copy of the catalog, which costs $.55. 970 customers made a purchase, spending an average of $75. The retailer's margin is 40%( variable costs are 60%)

a. what is the gross profit in dollars of this test mailing?

b. what is the gross profit % of gross sales?

c. What was the return on Marketing Expenditures (gross profit as a function of marketing costs?)

d. What was the breakeven
Answer #1

Expenses on catalogue = 0.55*30000 = 16500

Gross revenue = 970x75 = 72750

Margin for retailer = 0.4x72750= 29100

Gross profit = 29100-16500 = 12600

Profit as a percentage of sales = 100* ( 12600/72750) = 17.32%

Return on marketing expenditures = gross profit / marketing expenses = 12600/16500 = 0.7636

Breakeven is the point where profits are just good enough to meet costs

contribution per unit = 75x0.4 = 30

Number of customers needed to meet the fixed cost ( of marketing)

= 16500/30 = 550

which is the breakeven volume of sales